OpenStudy (anonymous):

What is the discriminant of 3x2 – 10x + 4 = 0?

OpenStudy (anonymous):

d = sqrt{b^2-4ac} b = -10 a = 3 c = 4 b^2 - 4ac = 100 - 48 d = sqrt{52}

OpenStudy (anonymous):

100-4*3*4 = 52